V=1/3PiR^2H how do you solve this problum if your solving for H?
Answer by nyc_function(2741)

V = (1/3)(pi)(R^2)H
V = (pi)(R^2)H
3V = (pi)(R^2)H
3V/[(pi)(R^2)] = H